A couple has three children. Find each probability:

 a . All boys

 b. All girls or all boys

 c. Exactly two boys or two girls

d. At least one child of each gender gril.

 


1
Expert's answer
2021-11-23T12:25:01-0500

Order is not important


"BBB, BBG, BGG, GGG"

a.


"P(BBB)=\dfrac{1}{4}=0.25"

b.


"P(BBB\ or \ GGG)=\dfrac{2}{4}=\dfrac{1}{2}=0.5"

c.


"P(BBG\ or\ BGG)=\dfrac{2}{4}=\dfrac{1}{2}=0.5"

d.

"P(at\ least\ 1 \ of \ each\ gender)=\dfrac{2}{4}=\dfrac{1}{2}=0.5"